# Refer to caption Refer to caption

### v1

@@ 4. Results
We generate all results with our default trajectory sampling scheme of 8 initial start images and 4 pre-defined trajectories. Depending on the collision detection, we generate between 1.5K-2.5K images per scene. The complete three-stage geneartion process takes on average 7h on a single RTX3090 GPU.
@@
We evaluate our generated 3D scenes both qualitatively and quantitatively. We calculate CLIP Score (CS) (Radford et al., 2021) and Inception Score (IS) (Salimans et al., 2016) on RGB renderings from novel perspectives beyond the panoramic/centered viewpoints of the respective scenes. Additionally, we conduct a user study and ask $\text{n}{=}64$ participants about the Perceptual Quality ($PQ$) and the 3D-Consistency ($3DC$). We present them with rendered videos that move throughout the whole 3D scene from novel (non-centered) perspectives and ask to score the scenes on a scale of $1{-}5$.
@@ 4.1. Qualitative Comparisons
”airy Hamptons beach house” Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
”melodic crystal music forest” Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
In Figure 5, we compare the quality of rendered images against baselines and we provide additional results of our method in Figure 4 and Figure 7. In both cases, we split the visualization into two parts. First, we visualize a scene overview located at the center of the scene to evaluate the quality and diversity of generated 3D scenes. Second, we render novel views that explore the scene, i.e., they are rotated around objects and moved away from the center. This way, showcases the ability of each method to fully navigate their generated scenes.
The panorama-based methods DreamScene360 (Zhou et al., 2024) and LayerPano3D (Yang et al., 2024b) create compelling $360^{\circ}$ views, but they do not hallucinate unobserved regions beyond the panoramic viewpoint. Thus, their novel views contain visible distortions and occlusions. Text2Room (Höllein et al., 2023) and WonderWorld (Yu et al., 2024a) create high-detail textures due to their iterative unprojection steps, but suffer from error accumulation, i.e., their object geometry appears stretched out due to erroneous depth estimation. FlexWorld (Chen et al., 2025b) creates complete 3D scenes with compelling hallucinations (e.g., the furnace). However, their trajectory-based generation does not cover large viewpoint changes, and as such novel views rotated around objects still exhibit missing and distorted geometry (e.g., behind the couch/table). SEVA (Zhou et al., 2025) uses a single, continuous trajectory to generate an entire 3D scene given a single image as input. This enables them to generate novel geometry, that matches the style of the input image. However, the generated video frames contain noticeable flickering artifacts (see supplementary material), which become visible in the 3DGS renderings as duplicated objects (e.g., multiple overlayed couches/tables) or floating artifacts (e.g., the bottom-left of the overview rendering).
@@ 4.2. Quantitative Comparisons
Table 1. Quantitative comparison. We report 2D metrics and user study results, including: Clip Score (CS), Inception Score (IS), Perceptual Quality (PQ), and 3D-Consistency  (3DC). We evaluate all methods with novel views rendered from perspectives beyond the scene center, i.e., we quantify their ability to explore the scene from arbitrary camera viewpoints. Our method creates scenes with the highest text-prompt similarity and is preferred in our user study.
@@
We present the quantitative results averaged across the novel view renderings of multiple scenes in  Table 1. Our approach achieves the highest CLIP score, since our generated geometry remains consistent towards the scene prompt even beyond panoramic perspectives. This highlights the advantage of our panorama initialization: the scene scaffold is already created from the beginning, which makes inpainting around it easier than for the iterative T2I and video diffusion baselines, that have to hallucinate the novel content completely from scratch. We obtain competitive Inception scores, which signals that our generated images are as sharp as baseline methods. Users give the highest ratings to our method, which highlights the quality of our generated 3D scenes under challenging, novel perspectives.
@@ 4.3. Ablations
We generate multiple small videos along pre-defined trajectories, that start from each of the 8 initial panoramic images. We ablate the importance of having 8 initial images by only providing 4 or 1 of these images as condition during video generation (but still run all 32 trajectory generations). We observe that scene diversity and sharpness drop noticeably, i.e., generating entire scene parts from scratch is harder than inpainting around our initial scene scaffold. Concretely, when providing only a single panoramic image, the video diffusion model needs to hallucinate entire new areas in a 3D-consistent fashion. This oftentimes leads to oversimplified outputs (if using a low guidance scale) or oversatured frames (if using a high guidance scale). While providing 4 images helps to retain large scenes, their 3D-consistency severely drops, as the individual trajectories are less well-conditioned. This results in blurriness and floating artifacts in the 3D scene reconstruction.
@@ 4.4. Limitations
Our method generates high-quality 3D scenes from text as input, that can be fully explored, i.e., we can interactively move into scene parts and rotate around objects without quality degradation. Nevertheless, there are remaining limitations (see supplementary material). Our video diffusion model sometimes produces high-frequency flickering artifacts across trajectories, which manifest in blurry renderings. Extending our 3D scene optimization with uncertainty (Goli et al., 2024) or robustness (Sabour et al., 2023) weighting schemes specifically tailored for these common problems of video generators may further detect and ignore such outliers. Furthermore, our method uses a set of pre-defined trajectories, which explore and expand the initial scene scaffold. While it proved robust towards a large variety of room and prompt types, it limits the layout of the generated scenes. Incorporating semantic layout as condition (Schult et al., 2024; Zhang et al., 2023b) into our iterative scene is an interesting future direction. Lastly, our approach generates 3D scenes with baked-in material and lighting effects, since our base diffusion models operate in RGB space. Recent text-to-x generators like (Kocsis et al., 2025; Ke et al., 2025) have the potential to replace RGB-space models in downstream tasks such as scene generation.
@@ Acknowledgements.
This project was funded by the ERC Consolidator Grant Gen3D (101171131) as well as the German Research Foundation (DFG) Research Unit “Learning and Simulation in Visual Computing”. We also thank Angela Dai for the video voice-over.
@@ Appendix A Supplemental Video
Please watch our attached video for a comprehensive evaluation of the proposed method. We include rendered videos of multiple generated scenes from novel trajectories, that showcase the quality of our large-scale 3D scenes under various non-centered perspectives. We also compare against baselines and ablations of our method by showing rendered videos.
@@ Appendix B User Study Details
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
We conduct a user study to determine Perceptual Quality (PQ) and 3D Consistency (3DC) of our 3D scenes. In order to determine these scores, we asked participants to rate our scenes and baselines based on overall quality and 3D-consistency on a scale of $1{-}5$. Figure 8 shows an example of how we asked the users to score the above two metrics. We show the users short videos of trajectories rendered from our final scene, as well as from baselines, both displaying large camera motion. In total, we received 2432 datapoints from $n{=}64$ participants.
@@ Appendix C Additional Baseline Details
Panorama-To-3D Generators
@@ Appendix D Limitations
Appendix E Additional Qualitative Results

### v2

@@ 4. Results
We generate all results with our default trajectory sampling scheme of 8 initial start images and 4 pre-defined trajectories. All 32 videos are generated sequentially in a clockwise order of starting images. Within each we first generate the “zoom-in” trajectory (1) followed by “left” (2), “right” (3), and “up” (4). We generate 44 frames per video for the trajectory types (1) and (4), and 134 frames for (2) and (3), respectively. Depending on the collision detection, we generate between 1.5K-2.5K images per scene at resolution $576\times 576$. The complete three-stage generation process takes on average 7h on a single RTX3090 GPU. Our runtime is dominated by the second stage (~10min/trajectory), while the first stage (~5min) and third stage (~11min) are comparatively faster.
@@
We evaluate our generated 3D scenes both qualitatively and quantitatively. We calculate CLIP Score (CS) (Radford et al., 2021) and Inception Score (IS) (Salimans et al., 2016) on RGB renderings from novel perspectives. For each scene, we sample cameras beyond panoramic/centered perspectives, to assess the quality under arbitrary viewpoints. Additionally, we conduct a user study and ask $\text{n}{=}64$ participants about the Perceptual Quality ($PQ$) and the 3D-Consistency ($3DC$). We present them with rendered videos that move throughout the whole 3D scene from novel (non-centered) perspectives and ask to score the scenes on a scale of 151{-}5.
@@ 4.1. Qualitative Comparisons
DreamScene360 (Zhou et al., 2024) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
LayerPano3D (Yang et al., 2024b) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Text2Room (Höllein et al., 2023) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
WonderWorld (Yu et al., 2024a) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
FlexWorld (Chen et al., 2025b) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SEVA (Zhou et al., 2025) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4. Qualitative comparison of our method and baselines. We generate a scene given the text prompt “star-written, constellation-binding, midnight celestial ink library”. We visualize a scene overview from a centered camera position and 4 novel views that explore the scene from non-centered perspectives. Our method retains compelling visual quality and generates plausible continuations of the scene, even under these extreme camera positions. Please see the supplementary material for animated comparisons against baselines.
[…]
Qualitative comparison of our method and baselines. We generate a scene given the text prompt “star-written, constellation-binding, midnight celestial ink library”. We visualize a scene overview from a centered camera position and 4 novel views that explore the scene from non-centered perspectives. Our method retains compelling visual quality and generates plausible continuations of the scene, even under these extreme camera positions. Please see the supplementary material for animated comparisons against baselines.
We compare the quality of generated scenes against all baselines in Figure 4. Also, we provide additional results of our method in Figure 6 and the supplementary material. In both cases, we split the visualization into two parts. First, we visualize a scene overview located at the center of the scene to evaluate the quality and diversity of generated 3D scenes. Second, we render novel views that explore the scene, i.e., they are rotated around objects and moved away from the center. This allows us to showcase the ability of each method to fully navigate their generated scenes.
The panorama-based methods DreamScene360 (Zhou et al., 2024) and LayerPano3D (Yang et al., 2024b) create compelling $360^{\circ}$ views, but they do not hallucinate unobserved regions beyond the panoramic viewpoint. Thus, their novel views contain visible distortions and occlusions. Text2Room (Höllein et al., 2023) and WonderWorld (Yu et al., 2024a) create high-detail textures due to their iterative unprojection steps, but suffer from error accumulation, i.e., their object geometry appears stretched out due to erroneous depth estimation. FlexWorld (Chen et al., 2025b) creates complete 3D scenes with compelling hallucinations (e.g., the furnace). However, their trajectory-based generation does not cover large viewpoint changes, and as such novel views rotated around objects still exhibit missing and distorted geometry (e.g., behind the couch/table). SEVA (Zhou et al., 2025) uses a single, continuous trajectory to generate an entire 3D scene given a single image as input. This enables them to generate novel geometry, that matches the style of the input image. However, the generated video frames contain noticeable flickering artifacts (see suppl. material), which become visible in the 3DGS renderings as duplicated objects (e.g., multiple overlayed couches/tables) or floating artifacts (e.g., the bottom-left of the overview rendering).
@@ 4.2. Quantitative Comparisons
Table 1. Quantitative comparison. We report 2D metrics and user study results, including: Clip Score (CS), Inception Score (IS), Perceptual Quality (PQ), and 3D-Consistency  (3DC). We evaluate all methods on novel views rendered from perspectives beyond the scene center, i.e., we quantify their ability to explore the scene arbitrarily. Our method creates scenes with the highest text-prompt similarity and is preferred in our user study.
@@
We present the quantitative results averaged across the novel view renderings of multiple scenes in  Table 1. Our approach achieves the highest CLIP score, since our generated geometry remains consistent towards the scene prompt even beyond panoramic perspectives. This highlights the advantage of our panorama initialization: the scene scaffold is already created from the beginning, which makes inpainting around it easier than for the iterative T2I and video diffusion baselines, that have to hallucinate the novel content completely from scratch. We obtain competitive Inception scores, which signals that our generated images are as sharp as baseline methods. Users give the highest ratings to our method, which highlights the quality of our generated 3D scenes under challenging, novel perspectives. We also provide a runtime and memory comparison against baselines in the supplementary material.
@@ 4.3. Ablations
Generated scene overview Rendered novel views
Ours 1 image Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours 4 images Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours w/o scene-mem Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours w/o coll-det Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5. Ablation study on the key components of our method. When conditioning video generation on only 1 initial image, generated scenes become smaller and less diverse: it is harder to generate novel areas from scratch than to inpaint around existing panoramic observations. While 4 initial panoramic images enable the creation of larger scenes, the frames of trajectories without a starting image are less 3D-consistent. This results in floating artifacts during rendering. Our proposed scene memory mechanism is necessary, as each individual trajectory only covers a local area of the scene. Thus, conditioning only within that trajectory similarly creates 3D-inconsistent scenes. Our collision detection ensures, that the pre-defined trajectories do not intersect with generated scene content like the archway towards the kitchen area. Without it, these areas become more fuzzy due to degenerate videos running into these structures.
[…]
Ablation study on the key components of our method. When conditioning video generation on only 1 initial image, generated scenes become smaller and less diverse: it is harder to generate novel areas from scratch than to inpaint around existing panoramic observations. While 4 initial panoramic images enable the creation of larger scenes, the frames of trajectories without a starting image are less 3D-consistent. This results in floating artifacts during rendering. Our proposed scene memory mechanism is necessary, as each individual trajectory only covers a local area of the scene. Thus, conditioning only within that trajectory similarly creates 3D-inconsistent scenes. Our collision detection ensures, that the pre-defined trajectories do not intersect with generated scene content like the archway towards the kitchen area. Without it, these areas become more fuzzy due to degenerate videos running into these structures.
@@
We generate multiple small videos along pre-defined trajectories, that start from each of the 8 initial panoramic images. We ablate the importance of having 8 initial images by only providing 4 or 1 of them as condition during video generation (but still generate all 32 trajectories). Scene diversity and sharpness drop noticeably, i.e., generating entire scene parts from scratch is harder than inpainting around our initial scene scaffold. Concretely, when providing only a single panoramic image, the video diffusion model needs to hallucinate entire new areas in a 3D-consistent fashion. This oftentimes leads to oversimplified outputs (if using a low guidance scale) or oversaturated frames (if using a high guidance scale). While providing 4 images helps to retain large scenes, their 3D-consistency severely drops, as the individual trajectories are less well-conditioned. This results in blurriness and floating artifacts in the 3D scene reconstruction.
@@ 4.4. Limitations
Generated scene overview Rendered novel views
”airy Hamptons beach house”
”melodic crystal music forest”
”liquid metal monastery”
[…]
World generation results of our method. We visualize scene overview renderings close to the scene center facing in two opposite directions. Additionally, we render novel views far beyond the scene center, that explore the different generated areas in more detail. This showcases the possibility to interactively explore our worlds from arbitrary viewpoints. We also demonstrate the diversity of our scenes by mixing indoor/outdoor elements in a single 3D scene. Please see the supplementary material for animated flythroughs of our generated scenes.
Our method generates high-quality 3D scenes from text as input, that can be fully explored, i.e., we can interactively move into scene parts and rotate around objects without quality degradation. Nevertheless, there are remaining limitations (see supplementary material). Our video diffusion model sometimes produces high-frequency flickering artifacts across trajectories, which manifest in blurry renderings. Extending our 3D scene optimization with uncertainty (Goli et al., 2024) or robustness (Sabour et al., 2023) weighting schemes specifically tailored for these common problems of video generators may further detect and ignore such outliers. Furthermore, our method uses a set of pre-defined trajectories, which explore and expand the initial scene scaffold. While it proved robust towards a large variety of room and prompt types, it limits the layout of the generated scenes. Incorporating semantic layouts as condition (Schult et al., 2024; Zhang et al., 2023a), enabling interactive and unlimited scene generations, and generating dynamic 4D worlds are interesting future directions. Lastly, our approach generates 3D scenes with baked-in material and lighting effects, since our base diffusion models operate in RGB space. Recent text-to-x generators like (Kocsis et al., 2025; Ke et al., 2025) have the potential to replace RGB-space models in downstream tasks such as scene generation.
@@ Acknowledgements.
This project was funded by the ERC Consolidator Grant Gen3D (101171131) as well as the German Research Foundation (DFG) Research Unit “Learning and Simulation in Visual Computing”. We also thank Angela Dai for the video voice-over and Simon Giebenhain for help on the Gaussian overlay visualization.
@@ Appendix A Supplemental Video
Please watch our attached video  [^1] https://youtu.be/N6NJsNyiv6I for a comprehensive evaluation of the proposed method. We include rendered videos of multiple generated scenes from novel trajectories, that showcase the quality of our large-scale 3D scenes under various non-centered perspectives. We also compare against baselines and ablations of our method by showing rendered videos.
@@ Appendix B User Study Details
[…]
User study interface. (Top) We provide users with videos along trajectories with large camera motion for our and baseline scenes. (Left) We ask users to scale the Perceptual Quality (PC) and 3D Consistency (3DC) on a scale from 1 to 5.
We conduct a user study to determine Perceptual Quality (PQ) and 3D Consistency (3DC) of our 3D scenes. In order to determine these scores, we asked participants to rate our scenes and baselines based on overall quality and 3D-consistency on a scale of 151{-}5. Figure 7 shows an example of how we asked the users to score the above two metrics. We show the users short videos of trajectories rendered from our final scene, as well as from baselines, both displaying large camera motion. In total, we received 2432 datapoints from n=64n{=}64 participants.
@@ Appendix C Additional Baseline Details
[…]
Intermediate baseline outputs. We compare the panorama generation of (Yang et al., 2024b) against our first stage panorama initialization (Section 3.2). Since we generate four images independently from each other, we can obtain environments with higher diversity than the baselines (e.g., it is possible to specify certain room types like “living room” or “kitchen” separately for each of the four images).
@@
[…]
Intermediate video generation results of  (Zhou et al., 2025). The above frames represents the video output generated along a single, continuous trajectory from the SEVA baseline (Zhou et al., 2025) initialized from a single input image on the “celestial ink library” scene. Even though we have chosen a high guidance scale of 6.0 for the generation, we still obtain inconsistent results (e.g. multiple different tables are generated in the bottom row). This leads to visible blurriness and duplicated object results when optimizing a 3DGS  (Kerbl et al., 2023) representation from these images. In contrast, our full pipeline divides scene generation into multiple smaller videos that are further conditioned on 8 initial panoramic images. The resulting scenes from our approach exhibit 3D-consistent and high quality renderings.
@@ Appendix D Limitations
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
[…]
Limitations of our method. Top and mid: high-frequency flickering in generated video frames lead to floating artifacts and blurry textures for certain parts of our 3D reconstruction. Bottom: objects placed at the very edge of our generated scene are rarely observed from all sides during video trajectory generation. Thus, the reconstructed scene is underconstrained at these areas, which results in degenerate rendering results.
@@
Appendix E Runtime/Memory Comparison
Table 2. Runtime/Memory comparison. Our method generates fully-explorable scenes, as opposed to baseline-scenes viewable from limited camera-poses. We generate more observations in order to resolve occlusions and increase the 3D completeness. The runtime amortized per generated frame on the “celestial library” scene is comparable to baselines.
Method Total runtime Generated frames Runtime per frame Memory
DreamScene360 (Zhou et al., 2024) 0.5h one panorama N/A 16GB
LayerPano3D (Yang et al., 2024b) 0.2h one panorama N/A 5GB
Text2Room (Höllein et al., 2023) 0.9h 200 16.2s 15GB
WonderWorld (Yu et al., 2024a) user-specific interactive generation ~10s 45GB
FlexWorld (Chen et al., 2025b) 0.5h 144 12.9s 48GB
SEVA (Zhou et al., 2025) 0.9h 270 13.5s 22GB
Ours 7h 2251 14.1s 22GB
Our method generates fully-explorable scenes, as opposed to baseline-scenes viewable from limited camera-poses. We generate more observations in order to resolve occlusions and increase the 3D completeness. This yields higher-quality scenes, as measured by Perceptual-Quality (PQ) and 3D-Consistency (3DC) of our user-study in Table 1. Therefore, we report runtime amortized per generated frame on the “celestial library” scene in Table 2. We achieve comparable per-frame generation rates and memory consumption as baselines approaches. Our runtime is dominated by the second stage (~10min/trajectory), while the first stage (~5min) and third stage (~11min) are comparatively faster. We believe real-time video-diffusion will reduce runtime towards interactive generation.
Appendix F Scene Layout Diversity
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
The 8 initial panoramic images corresponding to the prefixes: dining room $\rightarrow$ kitchen $\rightarrow$ living room $\rightarrow$ bedroom
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
The 8 initial panoramic images corresponding to the prefixes: wall $\rightarrow$ kitchen $\rightarrow$ living room $\rightarrow$ wall
Refer to caption Refer to caption Refer to caption
Scene rendering: kitchen $\rightarrow$ living room Scene rendering: dining room $\rightarrow$ kitchen Scene rendering: wall $\rightarrow$ kitchen
Figure 11. Scene layout diversity. We generate an ”urban industrial loft” with two different layouts by modifying the text prompts in the first stage. Top: generated panoramic scene scaffold with the room prefixes “dining room, kitchen, living room, bedroom”. Mid: replacing “dining room” and “bedroom” prefixes with “wall”. Bottom left: in both cases, we obtain scenes that transition from a kitchen to a living room. Bottom mid: in the first case (top) the generated scene contains another transition from a dining room area to the kitchen. Bottom right: in the second case (mid) we instead create a smaller scene without another room, but instead a flat wall transitioning to the kitchen.
[…]
Scene layout diversity. We generate an ”urban industrial loft” with two different layouts by modifying the text prompts in the first stage. Top: generated panoramic scene scaffold with the room prefixes “dining room, kitchen, living room, bedroom”. Mid: replacing “dining room” and “bedroom” prefixes with “wall”. Bottom left: in both cases, we obtain scenes that transition from a kitchen to a living room. Bottom mid: in the first case (top) the generated scene contains another transition from a dining room area to the kitchen. Bottom right: in the second case (mid) we instead create a smaller scene without another room, but instead a flat wall transitioning to the kitchen.
Our panoramic scaffold (Section 3.2) determines the global scene layout, i.e., which room types are combined. In our main results, we create the scaffold by using the same text prompt, but four different room prefixes (e.g. “kitchen/office of XYZ”). This creates scenes with a “cross-shaped” layout, where each of the four rooms becomes an explorable area.
We showcase in Figure 11, that our method can also generate other layouts (e.g., “L-shaped” scenes) by adjusting the panoramic scaffold as desired. Concretely, we replace two room prefixes with “wall/window of XYZ” to generate flat, non-explorable areas. This demonstrates the diversity of our approach: it is possible to customize the generated scene layouts by adjusting the panoramic scene scaffold accordingly.
Appendix G Additional Qualitative Results
@@
[…]
Additional world generation results of our method. We visualize scene overview renderings close to the scene center facing in two opposite directions. Additionally, we render novel views far beyond the scene center, that explore the different generated areas in more detail. This showcases the possibility to interactively explore our worlds from arbitrary viewpoints. Please see the supplementary video for animated flythroughs of our generated scenes.
@@
[…]
Additional world generation results of our method. We visualize scene overview renderings close to the scene center facing in two opposite directions. Additionally, we render novel views far beyond the scene center, that explore the different generated areas in more detail. This showcases the possibility to interactively explore our worlds from arbitrary viewpoints. Please see the supplementary video for animated flythroughs of our generated scenes.
@@
[…]
Additional world generation results of our method. We visualize scene overview renderings close to the scene center facing in two opposite directions. Additionally, we render novel views far beyond the scene center, that explore the different generated areas in more detail. This showcases the possibility to interactively explore our worlds from arbitrary viewpoints. Please see the supplementary video for animated flythroughs of our generated scenes.